An electron and a proton are lying 10 cm apart The ratio of electrostatics force and gravitational force between them will be

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

me =9.1 * 01-31

mp = 1.6 * 10-27

qe = -1.6 *10-19

qp =1.6 *10-19

Fg =Gmemp/(10)2

Fe =1/4∏ε qeqp/(10)2

  Fe/Fg = (1/4∏ε) qeqp /Gmemp

             = 9*109 *(1.6 *10-19)2/(6.67*10-11 * 9.1 * 10-31 *1.6 * 10-27 )

           =.23 * 10 40

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin